# A causal inference framework for spatial confounding

### v7

@@ 8 Discussion
n method bias sd mse coverage
$1.000\times 10^{3}$ RSR $-2.791\times 10^{0}$ $1.065\times 10^{0}$ $8.921\times 10^{0}$ 0%
$1.000\times 10^{3}$ spline $-1.157\times 10^{0}$ $6.826\times 10^{-1}$ $1.804\times 10^{0}$ 25%
$1.000\times 10^{3}$ gp $-1.322\times 10^{0}$ $6.891\times 10^{-1}$ $2.221\times 10^{0}$ 17%
$1.000\times 10^{3}$ gSEM $-1.254\times 10^{0}$ $7.258\times 10^{-1}$ $2.098\times 10^{0}$ 55%
$1.000\times 10^{3}$ spatial+ $-1.087\times 10^{0}$ $6.927\times 10^{-1}$ $1.661\times 10^{0}$ 63%
$1.000\times 10^{3}$ svc_mle $-1.145\times 10^{0}$ $2.002\times 10^{-1}$ $1.350\times 10^{0}$
$1.000\times 10^{3}$ spline_interaction $-9.852\times 10^{-1}$ $4.471\times 10^{-1}$ $1.170\times 10^{0}$ 38%
$1.000\times 10^{3}$ gp3 $3.652\times 10^{-2}$ $3.773\times 10^{-1}$ $1.434\times 10^{-1}$
$1.000\times 10^{3}$ spline3 $8.767\times 10^{-2}$ $3.791\times 10^{-1}$ $1.511\times 10^{-1}$
$1.000\times 10^{3}$ rf $-3.678\times 10^{-1}$ $3.409\times 10^{-1}$ $2.513\times 10^{-1}$ 75%
$1.000\times 10^{3}$ BART $2.818\times 10^{-2}$ $3.902\times 10^{-1}$ $1.527\times 10^{-1}$ 98%
$1.000\times 10^{3}$ DML_gp3 $2.911\times 10^{-2}$ $3.771\times 10^{-1}$ $\mathbf{1.428\times 10^{-1}}$
$1.000\times 10^{3}$ DML_spline $5.456\times 10^{-2}$ $3.765\times 10^{-1}$ $1.444\times 10^{-1}$
$1.000\times 10^{3}$ DML_BART $\mathbf{4.759\times 10^{-3}}$ $3.901\times 10^{-1}$ $1.519\times 10^{-1}$ 98%
$1.000\times 10^{4}$ RSR $-2.853\times 10^{0}$ $3.774\times 10^{-1}$ $8.283\times 10^{0}$ 0%
$1.000\times 10^{4}$ spline $-1.061\times 10^{0}$ $2.014\times 10^{-1}$ $1.166\times 10^{0}$ 0%
$1.000\times 10^{4}$ gp $-1.084\times 10^{0}$ $2.010\times 10^{-1}$ $1.215\times 10^{0}$ 0%
$1.000\times 10^{4}$ gSEM $-1.146\times 10^{0}$ $2.053\times 10^{-1}$ $1.356\times 10^{0}$ 0%
$1.000\times 10^{4}$ spatial+ $-1.042\times 10^{0}$ $2.014\times 10^{-1}$ $1.126\times 10^{0}$ 0%
$1.000\times 10^{4}$ spline_interaction $-9.893\times 10^{-1}$ $1.594\times 10^{-1}$ $1.004\times 10^{0}$ 0%
$1.000\times 10^{4}$ spline3 $2.979\times 10^{-2}$ $1.229\times 10^{-1}$ $1.594\times 10^{-2}$
$1.000\times 10^{4}$ rf $-1.456\times 10^{-1}$ $1.143\times 10^{-1}$ $3.421\times 10^{-2}$ 75%
$1.000\times 10^{4}$ BART $1.308\times 10^{-2}$ $1.253\times 10^{-1}$ $1.581\times 10^{-2}$ 96%
$1.000\times 10^{4}$ DML_spline $1.421\times 10^{-2}$ $1.229\times 10^{-1}$ $\mathbf{1.524\times 10^{-2}}$
$1.000\times 10^{4}$ DML_BART $\mathbf{5.706\times 10^{-3}}$ $1.254\times 10^{-1}$ $1.569\times 10^{-2}$ 96%
@@
n method bias sd mse coverage
$1.000\times 10^{3}$ RSR $1.731\times 10^{-1}$ $8.450\times 10^{-2}$ $3.707\times 10^{-2}$ 0%
$1.000\times 10^{3}$ spline $2.433\times 10^{-2}$ $8.790\times 10^{-2}$ $8.302\times 10^{-3}$ 51%
$1.000\times 10^{3}$ gp $3.795\times 10^{-2}$ $8.729\times 10^{-2}$ $9.044\times 10^{-3}$ 51%
$1.000\times 10^{3}$ gSEM $5.425\times 10^{-2}$ $9.213\times 10^{-2}$ $1.141\times 10^{-2}$ 90%
$1.000\times 10^{3}$ spatial+ $2.323\times 10^{-2}$ $9.081\times 10^{-2}$ $8.770\times 10^{-3}$ 93%
$1.000\times 10^{3}$ svc_mle $2.079\times 10^{-2}$ $3.639\times 10^{-2}$ $\mathbf{1.754\times 10^{-3}}$
$1.000\times 10^{3}$ spline_interaction $4.651\times 10^{-2}$ $5.233\times 10^{-2}$ $4.896\times 10^{-3}$ 94%
$1.000\times 10^{3}$ gp3 $1.817\times 10^{-2}$ $1.390\times 10^{-1}$ $1.962\times 10^{-2}$
$1.000\times 10^{3}$ spline3 $4.436\times 10^{-2}$ $2.268\times 10^{-1}$ $5.331\times 10^{-2}$
$1.000\times 10^{3}$ rf $-1.103\times 10^{-1}$ $5.951\times 10^{-2}$ $1.571\times 10^{-2}$ 72%
$1.000\times 10^{3}$ BART $1.216\times 10^{-2}$ $1.014\times 10^{-1}$ $1.040\times 10^{-2}$ 100%
$1.000\times 10^{3}$ DML_gp3 $-1.420\times 10^{-2}$ $1.363\times 10^{-1}$ $1.874\times 10^{-2}$
$1.000\times 10^{3}$ DML_spline $2.872\times 10^{-2}$ $2.301\times 10^{-1}$ $5.369\times 10^{-2}$
$1.000\times 10^{3}$ DML_BART $\mathbf{6.093\times 10^{-3}}$ $1.023\times 10^{-1}$ $1.048\times 10^{-2}$ 100%
$1.000\times 10^{4}$ RSR $1.760\times 10^{-1}$ $2.928\times 10^{-2}$ $3.184\times 10^{-2}$ 0%
$1.000\times 10^{4}$ spline $1.043\times 10^{-2}$ $2.919\times 10^{-2}$ $9.573\times 10^{-4}$ 52%
$1.000\times 10^{4}$ gp $2.034\times 10^{-2}$ $2.946\times 10^{-2}$ $1.278\times 10^{-3}$ 47%
$1.000\times 10^{4}$ gSEM $1.407\times 10^{-2}$ $2.939\times 10^{-2}$ $1.058\times 10^{-3}$ 90%
$1.000\times 10^{4}$ spatial+ $5.516\times 10^{-3}$ $2.923\times 10^{-2}$ $8.816\times 10^{-4}$ 93%
$1.000\times 10^{4}$ spline_interaction $2.210\times 10^{-2}$ $1.800\times 10^{-2}$ $\mathbf{8.110\times 10^{-4}}$ 80%
$1.000\times 10^{4}$ spline3 $4.608\times 10^{-2}$ $2.875\times 10^{-2}$ $2.947\times 10^{-3}$
$1.000\times 10^{4}$ rf $-5.237\times 10^{-2}$ $2.216\times 10^{-2}$ $3.231\times 10^{-3}$ 62%
$1.000\times 10^{4}$ BART $1.036\times 10^{-2}$ $3.330\times 10^{-2}$ $1.212\times 10^{-3}$ 100%
$1.000\times 10^{4}$ DML_spline $1.740\times 10^{-2}$ $3.327\times 10^{-2}$ $1.405\times 10^{-3}$
$1.000\times 10^{4}$ DML_BART $\mathbf{1.913\times 10^{-3}}$ $3.456\times 10^{-2}$ $1.193\times 10^{-3}$ 100%
@@
n method bias sd mse coverage
$1.000\times 10^{3}$ RSR $4.240\times 10^{-1}$ $3.038\times 10^{-2}$ $1.807\times 10^{-1}$ 0%
$1.000\times 10^{3}$ spline $2.979\times 10^{-1}$ $2.432\times 10^{-2}$ $8.936\times 10^{-2}$ 0%
$1.000\times 10^{3}$ gp $3.004\times 10^{-1}$ $2.442\times 10^{-2}$ $9.086\times 10^{-2}$ 0%
$1.000\times 10^{3}$ gSEM $3.521\times 10^{-1}$ $2.521\times 10^{-2}$ $1.246\times 10^{-1}$ 0%
$1.000\times 10^{3}$ spatial+ $3.055\times 10^{-1}$ $2.443\times 10^{-2}$ $9.391\times 10^{-2}$ 0%
$1.000\times 10^{3}$ svc_mle $3.397\times 10^{-1}$ $2.318\times 10^{-2}$ $1.160\times 10^{-1}$
$1.000\times 10^{3}$ spline_interaction $3.436\times 10^{-1}$ $1.890\times 10^{-2}$ $1.184\times 10^{-1}$ 0%
$1.000\times 10^{3}$ gp3 $3.753\times 10^{-1}$ $2.096\times 10^{-2}$ $1.413\times 10^{-1}$
$1.000\times 10^{3}$ spline3 $4.230\times 10^{-1}$ $2.279\times 10^{-2}$ $1.795\times 10^{-1}$
$1.000\times 10^{3}$ rf $\mathbf{1.072\times 10^{-1}}$ $3.980\times 10^{-2}$ $\mathbf{1.308\times 10^{-2}}$ 49%
$1.000\times 10^{3}$ BART $3.278\times 10^{-1}$ $5.504\times 10^{-2}$ $1.105\times 10^{-1}$ 0%
$1.000\times 10^{3}$ DML_gp3 $3.572\times 10^{-1}$ $2.244\times 10^{-2}$ $1.281\times 10^{-1}$
$1.000\times 10^{3}$ DML_spline $3.324\times 10^{-1}$ $2.567\times 10^{-2}$ $1.111\times 10^{-1}$
$1.000\times 10^{3}$ DML_BART $3.204\times 10^{-1}$ $5.549\times 10^{-2}$ $1.057\times 10^{-1}$ 0%
$1.000\times 10^{4}$ RSR $4.197\times 10^{-1}$ $1.015\times 10^{-2}$ $1.763\times 10^{-1}$ 0%
$1.000\times 10^{4}$ spline $2.879\times 10^{-1}$ $7.850\times 10^{-3}$ $8.295\times 10^{-2}$ 0%
$1.000\times 10^{4}$ gp $2.880\times 10^{-1}$ $7.820\times 10^{-3}$ $8.299\times 10^{-2}$ 0%
$1.000\times 10^{4}$ gSEM $3.007\times 10^{-1}$ $8.043\times 10^{-3}$ $9.051\times 10^{-2}$ 0%
$1.000\times 10^{4}$ spatial+ $2.879\times 10^{-1}$ $7.798\times 10^{-3}$ $8.296\times 10^{-2}$ 0%
$1.000\times 10^{4}$ spline_interaction $3.443\times 10^{-1}$ $5.832\times 10^{-3}$ $1.186\times 10^{-1}$ 0%
$1.000\times 10^{4}$ spline3 $3.678\times 10^{-1}$ $1.214\times 10^{-2}$ $1.354\times 10^{-1}$
$1.000\times 10^{4}$ rf $\mathbf{2.182\times 10^{-1}}$ $1.239\times 10^{-2}$ $\mathbf{4.775\times 10^{-2}}$ 0%
$1.000\times 10^{4}$ BART $3.194\times 10^{-1}$ $2.167\times 10^{-2}$ $1.025\times 10^{-1}$ 0%
$1.000\times 10^{4}$ DML_spline $3.373\times 10^{-1}$ $1.202\times 10^{-2}$ $1.139\times 10^{-1}$
$1.000\times 10^{4}$ DML_BART $3.255\times 10^{-1}$ $2.052\times 10^{-2}$ $1.064\times 10^{-1}$ 0%
@@
n method bias sd mse coverage
$1.000\times 10^{3}$ RSR $4.338\times 10^{-1}$ $2.410\times 10^{-2}$ $1.888\times 10^{-1}$ 0%
$1.000\times 10^{3}$ spline $1.421\times 10^{-2}$ $7.492\times 10^{-3}$ $2.581\times 10^{-4}$ 49%
$1.000\times 10^{3}$ gp $1.464\times 10^{-2}$ $7.791\times 10^{-3}$ $2.748\times 10^{-4}$ 49%
$1.000\times 10^{3}$ gSEM $9.108\times 10^{-2}$ $9.040\times 10^{-3}$ $8.378\times 10^{-3}$ 0%
$1.000\times 10^{3}$ spatial+ $1.768\times 10^{-2}$ $7.902\times 10^{-3}$ $3.750\times 10^{-4}$ 45%
$1.000\times 10^{3}$ svc_mle $1.401\times 10^{-2}$ $7.693\times 10^{-3}$ $2.553\times 10^{-4}$
$1.000\times 10^{3}$ spline_interaction $1.403\times 10^{-2}$ $7.515\times 10^{-3}$ $2.532\times 10^{-4}$ 81%
$1.000\times 10^{3}$ gp3 $8.156\times 10^{-3}$ $7.658\times 10^{-3}$ $1.251\times 10^{-4}$
$1.000\times 10^{3}$ spline3 $9.351\times 10^{-2}$ $6.410\times 10^{-2}$ $1.284\times 10^{-2}$
$1.000\times 10^{3}$ rf $-1.034\times 10^{-1}$ $2.001\times 10^{-2}$ $1.109\times 10^{-2}$ 2%
$1.000\times 10^{3}$ BART $1.608\times 10^{-2}$ $1.896\times 10^{-2}$ $6.173\times 10^{-4}$ 100%
$1.000\times 10^{3}$ DML_gp3 $\mathbf{7.789\times 10^{-3}}$ $7.932\times 10^{-3}$ $\mathbf{1.235\times 10^{-4}}$
$1.000\times 10^{3}$ DML_spline $5.611\times 10^{-2}$ $5.497\times 10^{-2}$ $6.164\times 10^{-3}$
$1.000\times 10^{3}$ DML_BART $1.143\times 10^{-2}$ $1.889\times 10^{-2}$ $4.865\times 10^{-4}$ 100%
$1.000\times 10^{4}$ RSR $4.334\times 10^{-1}$ $7.335\times 10^{-3}$ $1.879\times 10^{-1}$ 0%
$1.000\times 10^{4}$ spline $8.398\times 10^{-3}$ $2.300\times 10^{-3}$ $7.579\times 10^{-5}$ 6%
$1.000\times 10^{4}$ gp $9.312\times 10^{-3}$ $2.298\times 10^{-3}$ $9.197\times 10^{-5}$ 2%
$1.000\times 10^{4}$ gSEM $2.439\times 10^{-2}$ $2.461\times 10^{-3}$ $6.009\times 10^{-4}$ 0%
$1.000\times 10^{4}$ spatial+ $7.380\times 10^{-3}$ $2.299\times 10^{-3}$ $5.973\times 10^{-5}$ 11%
$1.000\times 10^{4}$ spline_interaction $8.209\times 10^{-3}$ $2.306\times 10^{-3}$ $7.269\times 10^{-5}$ 7%
$1.000\times 10^{4}$ spline3 $4.054\times 10^{-2}$ $1.026\times 10^{-2}$ $1.748\times 10^{-3}$
$1.000\times 10^{4}$ rf $-5.936\times 10^{-2}$ $5.370\times 10^{-3}$ $3.553\times 10^{-3}$ 0%
$1.000\times 10^{4}$ BART $8.915\times 10^{-3}$ $3.818\times 10^{-3}$ $9.399\times 10^{-5}$ 64%
$1.000\times 10^{4}$ DML_spline $1.824\times 10^{-2}$ $1.001\times 10^{-2}$ $4.325\times 10^{-4}$
$1.000\times 10^{4}$ DML_BART $\mathbf{6.480\times 10^{-3}}$ $3.778\times 10^{-3}$ $\mathbf{5.621\times 10^{-5}}$ 84%

### v8

@@ 8 Discussion
{adjustwidth}
-3cm-3cm n method bias sd mse coverage $1.000\times 10^{3}$ RSR $-2.791\times 10^{0}$ $1.065\times 10^{0}$ $8.921\times 10^{0}$ 0% $1.000\times 10^{3}$ spline $-1.157\times 10^{0}$ $6.826\times 10^{-1}$ $1.804\times 10^{0}$ 25% $1.000\times 10^{3}$ gp $-1.322\times 10^{0}$ $6.891\times 10^{-1}$ $2.221\times 10^{0}$ 17% $1.000\times 10^{3}$ gSEM $-1.254\times 10^{0}$ $7.258\times 10^{-1}$ $2.098\times 10^{0}$ 55% $1.000\times 10^{3}$ spatial+ $-1.087\times 10^{0}$ $6.927\times 10^{-1}$ $1.661\times 10^{0}$ 63% $1.000\times 10^{3}$ svc_mle $-1.145\times 10^{0}$ $2.002\times 10^{-1}$ $1.350\times 10^{0}$ $1.000\times 10^{3}$ spline_interaction $-9.852\times 10^{-1}$ $4.471\times 10^{-1}$ $1.170\times 10^{0}$ 38% $1.000\times 10^{3}$ gp3 $3.652\times 10^{-2}$ $3.773\times 10^{-1}$ $1.434\times 10^{-1}$ $1.000\times 10^{3}$ spline3 $8.767\times 10^{-2}$ $3.791\times 10^{-1}$ $1.511\times 10^{-1}$ $1.000\times 10^{3}$ rf $-3.678\times 10^{-1}$ $3.409\times 10^{-1}$ $2.513\times 10^{-1}$ 75% $1.000\times 10^{3}$ BART $2.818\times 10^{-2}$ $3.902\times 10^{-1}$ $1.527\times 10^{-1}$ 98% $1.000\times 10^{3}$ DML_gp3 $2.911\times 10^{-2}$ $3.771\times 10^{-1}$ $\mathbf{1.428\times 10^{-1}}$ $1.000\times 10^{3}$ DML_spline $5.456\times 10^{-2}$ $3.765\times 10^{-1}$ $1.444\times 10^{-1}$ $1.000\times 10^{3}$ DML_BART $\mathbf{4.759\times 10^{-3}}$ $3.901\times 10^{-1}$ $1.519\times 10^{-1}$ 98% $1.000\times 10^{4}$ RSR $-2.853\times 10^{0}$ $3.774\times 10^{-1}$ $8.283\times 10^{0}$ 0% $1.000\times 10^{4}$ spline $-1.061\times 10^{0}$ $2.014\times 10^{-1}$ $1.166\times 10^{0}$ 0% $1.000\times 10^{4}$ gp $-1.084\times 10^{0}$ $2.010\times 10^{-1}$ $1.215\times 10^{0}$ 0% $1.000\times 10^{4}$ gSEM $-1.146\times 10^{0}$ $2.053\times 10^{-1}$ $1.356\times 10^{0}$ 0% $1.000\times 10^{4}$ spatial+ $-1.042\times 10^{0}$ $2.014\times 10^{-1}$ $1.126\times 10^{0}$ 0% $1.000\times 10^{4}$ spline_interaction $-9.893\times 10^{-1}$ $1.594\times 10^{-1}$ $1.004\times 10^{0}$ 0% $1.000\times 10^{4}$ spline3 $2.979\times 10^{-2}$ $1.229\times 10^{-1}$ $1.594\times 10^{-2}$ $1.000\times 10^{4}$ rf $-1.456\times 10^{-1}$ $1.143\times 10^{-1}$ $3.421\times 10^{-2}$ 75% $1.000\times 10^{4}$ BART $1.308\times 10^{-2}$ $1.253\times 10^{-1}$ $1.581\times 10^{-2}$ 96% $1.000\times 10^{4}$ DML_spline $1.421\times 10^{-2}$ $1.229\times 10^{-1}$ $\mathbf{1.524\times 10^{-2}}$ $1.000\times 10^{4}$ DML_BART $\mathbf{5.706\times 10^{-3}}$ $1.254\times 10^{-1}$ $1.569\times 10^{-2}$ 96%
@@
{adjustwidth}
-3cm-3cm n method bias sd mse coverage $1.000\times 10^{3}$ RSR $1.731\times 10^{-1}$ $8.450\times 10^{-2}$ $3.707\times 10^{-2}$ 0% $1.000\times 10^{3}$ spline $2.433\times 10^{-2}$ $8.790\times 10^{-2}$ $8.302\times 10^{-3}$ 51% $1.000\times 10^{3}$ gp $3.795\times 10^{-2}$ $8.729\times 10^{-2}$ $9.044\times 10^{-3}$ 51% $1.000\times 10^{3}$ gSEM $5.425\times 10^{-2}$ $9.213\times 10^{-2}$ $1.141\times 10^{-2}$ 90% $1.000\times 10^{3}$ spatial+ $2.323\times 10^{-2}$ $9.081\times 10^{-2}$ $8.770\times 10^{-3}$ 93% $1.000\times 10^{3}$ svc_mle $2.079\times 10^{-2}$ $3.639\times 10^{-2}$ $\mathbf{1.754\times 10^{-3}}$ $1.000\times 10^{3}$ spline_interaction $4.651\times 10^{-2}$ $5.233\times 10^{-2}$ $4.896\times 10^{-3}$ 94% $1.000\times 10^{3}$ gp3 $1.817\times 10^{-2}$ $1.390\times 10^{-1}$ $1.962\times 10^{-2}$ $1.000\times 10^{3}$ spline3 $4.436\times 10^{-2}$ $2.268\times 10^{-1}$ $5.331\times 10^{-2}$ $1.000\times 10^{3}$ rf $-1.103\times 10^{-1}$ $5.951\times 10^{-2}$ $1.571\times 10^{-2}$ 72% $1.000\times 10^{3}$ BART $1.216\times 10^{-2}$ $1.014\times 10^{-1}$ $1.040\times 10^{-2}$ 100% $1.000\times 10^{3}$ DML_gp3 $-1.420\times 10^{-2}$ $1.363\times 10^{-1}$ $1.874\times 10^{-2}$ $1.000\times 10^{3}$ DML_spline $2.872\times 10^{-2}$ $2.301\times 10^{-1}$ $5.369\times 10^{-2}$ $1.000\times 10^{3}$ DML_BART $\mathbf{6.093\times 10^{-3}}$ $1.023\times 10^{-1}$ $1.048\times 10^{-2}$ 100% $1.000\times 10^{4}$ RSR $1.760\times 10^{-1}$ $2.928\times 10^{-2}$ $3.184\times 10^{-2}$ 0% $1.000\times 10^{4}$ spline $1.043\times 10^{-2}$ $2.919\times 10^{-2}$ $9.573\times 10^{-4}$ 52% $1.000\times 10^{4}$ gp $2.034\times 10^{-2}$ $2.946\times 10^{-2}$ $1.278\times 10^{-3}$ 47% $1.000\times 10^{4}$ gSEM $1.407\times 10^{-2}$ $2.939\times 10^{-2}$ $1.058\times 10^{-3}$ 90% $1.000\times 10^{4}$ spatial+ $5.516\times 10^{-3}$ $2.923\times 10^{-2}$ $8.816\times 10^{-4}$ 93% $1.000\times 10^{4}$ spline_interaction $2.210\times 10^{-2}$ $1.800\times 10^{-2}$ $\mathbf{8.110\times 10^{-4}}$ 80% $1.000\times 10^{4}$ spline3 $4.608\times 10^{-2}$ $2.875\times 10^{-2}$ $2.947\times 10^{-3}$ $1.000\times 10^{4}$ rf $-5.237\times 10^{-2}$ $2.216\times 10^{-2}$ $3.231\times 10^{-3}$ 62% $1.000\times 10^{4}$ BART $1.036\times 10^{-2}$ $3.330\times 10^{-2}$ $1.212\times 10^{-3}$ 100% $1.000\times 10^{4}$ DML_spline $1.740\times 10^{-2}$ $3.327\times 10^{-2}$ $1.405\times 10^{-3}$ $1.000\times 10^{4}$ DML_BART $\mathbf{1.913\times 10^{-3}}$ $3.456\times 10^{-2}$ $1.193\times 10^{-3}$ 100%
@@
{adjustwidth}
-3cm-3cm n method bias sd mse coverage $1.000\times 10^{3}$ RSR $4.240\times 10^{-1}$ $3.038\times 10^{-2}$ $1.807\times 10^{-1}$ 0% $1.000\times 10^{3}$ spline $2.979\times 10^{-1}$ $2.432\times 10^{-2}$ $8.936\times 10^{-2}$ 0% $1.000\times 10^{3}$ gp $3.004\times 10^{-1}$ $2.442\times 10^{-2}$ $9.086\times 10^{-2}$ 0% $1.000\times 10^{3}$ gSEM $3.521\times 10^{-1}$ $2.521\times 10^{-2}$ $1.246\times 10^{-1}$ 0% $1.000\times 10^{3}$ spatial+ $3.055\times 10^{-1}$ $2.443\times 10^{-2}$ $9.391\times 10^{-2}$ 0% $1.000\times 10^{3}$ svc_mle $3.397\times 10^{-1}$ $2.318\times 10^{-2}$ $1.160\times 10^{-1}$ $1.000\times 10^{3}$ spline_interaction $3.436\times 10^{-1}$ $1.890\times 10^{-2}$ $1.184\times 10^{-1}$ 0% $1.000\times 10^{3}$ gp3 $3.753\times 10^{-1}$ $2.096\times 10^{-2}$ $1.413\times 10^{-1}$ $1.000\times 10^{3}$ spline3 $4.230\times 10^{-1}$ $2.279\times 10^{-2}$ $1.795\times 10^{-1}$ $1.000\times 10^{3}$ rf $\mathbf{1.072\times 10^{-1}}$ $3.980\times 10^{-2}$ $\mathbf{1.308\times 10^{-2}}$ 49% $1.000\times 10^{3}$ BART $3.278\times 10^{-1}$ $5.504\times 10^{-2}$ $1.105\times 10^{-1}$ 0% $1.000\times 10^{3}$ DML_gp3 $3.572\times 10^{-1}$ $2.244\times 10^{-2}$ $1.281\times 10^{-1}$ $1.000\times 10^{3}$ DML_spline $3.324\times 10^{-1}$ $2.567\times 10^{-2}$ $1.111\times 10^{-1}$ $1.000\times 10^{3}$ DML_BART $3.204\times 10^{-1}$ $5.549\times 10^{-2}$ $1.057\times 10^{-1}$ 0% $1.000\times 10^{4}$ RSR $4.197\times 10^{-1}$ $1.015\times 10^{-2}$ $1.763\times 10^{-1}$ 0% $1.000\times 10^{4}$ spline $2.879\times 10^{-1}$ $7.850\times 10^{-3}$ $8.295\times 10^{-2}$ 0% $1.000\times 10^{4}$ gp $2.880\times 10^{-1}$ $7.820\times 10^{-3}$ $8.299\times 10^{-2}$ 0% $1.000\times 10^{4}$ gSEM $3.007\times 10^{-1}$ $8.043\times 10^{-3}$ $9.051\times 10^{-2}$ 0% $1.000\times 10^{4}$ spatial+ $2.879\times 10^{-1}$ $7.798\times 10^{-3}$ $8.296\times 10^{-2}$ 0% $1.000\times 10^{4}$ spline_interaction $3.443\times 10^{-1}$ $5.832\times 10^{-3}$ $1.186\times 10^{-1}$ 0% $1.000\times 10^{4}$ spline3 $3.678\times 10^{-1}$ $1.214\times 10^{-2}$ $1.354\times 10^{-1}$ $1.000\times 10^{4}$ rf $\mathbf{2.182\times 10^{-1}}$ $1.239\times 10^{-2}$ $\mathbf{4.775\times 10^{-2}}$ 0% $1.000\times 10^{4}$ BART $3.194\times 10^{-1}$ $2.167\times 10^{-2}$ $1.025\times 10^{-1}$ 0% $1.000\times 10^{4}$ DML_spline $3.373\times 10^{-1}$ $1.202\times 10^{-2}$ $1.139\times 10^{-1}$ $1.000\times 10^{4}$ DML_BART $3.255\times 10^{-1}$ $2.052\times 10^{-2}$ $1.064\times 10^{-1}$ 0%
@@
{adjustwidth}
-3cm-3cm n method bias sd mse coverage $1.000\times 10^{3}$ RSR $4.338\times 10^{-1}$ $2.410\times 10^{-2}$ $1.888\times 10^{-1}$ 0% $1.000\times 10^{3}$ spline $1.421\times 10^{-2}$ $7.492\times 10^{-3}$ $2.581\times 10^{-4}$ 49% $1.000\times 10^{3}$ gp $1.464\times 10^{-2}$ $7.791\times 10^{-3}$ $2.748\times 10^{-4}$ 49% $1.000\times 10^{3}$ gSEM $9.108\times 10^{-2}$ $9.040\times 10^{-3}$ $8.378\times 10^{-3}$ 0% $1.000\times 10^{3}$ spatial+ $1.768\times 10^{-2}$ $7.902\times 10^{-3}$ $3.750\times 10^{-4}$ 45% $1.000\times 10^{3}$ svc_mle $1.401\times 10^{-2}$ $7.693\times 10^{-3}$ $2.553\times 10^{-4}$ $1.000\times 10^{3}$ spline_interaction $1.403\times 10^{-2}$ $7.515\times 10^{-3}$ $2.532\times 10^{-4}$ 81% $1.000\times 10^{3}$ gp3 $8.156\times 10^{-3}$ $7.658\times 10^{-3}$ $1.251\times 10^{-4}$ $1.000\times 10^{3}$ spline3 $9.351\times 10^{-2}$ $6.410\times 10^{-2}$ $1.284\times 10^{-2}$ $1.000\times 10^{3}$ rf $-1.034\times 10^{-1}$ $2.001\times 10^{-2}$ $1.109\times 10^{-2}$ 2% $1.000\times 10^{3}$ BART $1.608\times 10^{-2}$ $1.896\times 10^{-2}$ $6.173\times 10^{-4}$ 100% $1.000\times 10^{3}$ DML_gp3 $\mathbf{7.789\times 10^{-3}}$ $7.932\times 10^{-3}$ $\mathbf{1.235\times 10^{-4}}$ $1.000\times 10^{3}$ DML_spline $5.611\times 10^{-2}$ $5.497\times 10^{-2}$ $6.164\times 10^{-3}$ $1.000\times 10^{3}$ DML_BART $1.143\times 10^{-2}$ $1.889\times 10^{-2}$ $4.865\times 10^{-4}$ 100% $1.000\times 10^{4}$ RSR $4.334\times 10^{-1}$ $7.335\times 10^{-3}$ $1.879\times 10^{-1}$ 0% $1.000\times 10^{4}$ spline $8.398\times 10^{-3}$ $2.300\times 10^{-3}$ $7.579\times 10^{-5}$ 6% $1.000\times 10^{4}$ gp $9.312\times 10^{-3}$ $2.298\times 10^{-3}$ $9.197\times 10^{-5}$ 2% $1.000\times 10^{4}$ gSEM $2.439\times 10^{-2}$ $2.461\times 10^{-3}$ $6.009\times 10^{-4}$ 0% $1.000\times 10^{4}$ spatial+ $7.380\times 10^{-3}$ $2.299\times 10^{-3}$ $5.973\times 10^{-5}$ 11% $1.000\times 10^{4}$ spline_interaction $8.209\times 10^{-3}$ $2.306\times 10^{-3}$ $7.269\times 10^{-5}$ 7% $1.000\times 10^{4}$ spline3 $4.054\times 10^{-2}$ $1.026\times 10^{-2}$ $1.748\times 10^{-3}$ $1.000\times 10^{4}$ rf $-5.936\times 10^{-2}$ $5.370\times 10^{-3}$ $3.553\times 10^{-3}$ 0% $1.000\times 10^{4}$ BART $8.915\times 10^{-3}$ $3.818\times 10^{-3}$ $9.399\times 10^{-5}$ 64% $1.000\times 10^{4}$ DML_spline $1.824\times 10^{-2}$ $1.001\times 10^{-2}$ $4.325\times 10^{-4}$ $1.000\times 10^{4}$ DML_BART $\mathbf{6.480\times 10^{-3}}$ $3.778\times 10^{-3}$ $\mathbf{5.621\times 10^{-5}}$ 84%